OpenStudy (anonymous):

What is the inverse tan of tan((4*pi)/3)? Show your steps. I would have answered (4*pi)/3 but I wonder why the answer is just pi/3? Where did the four go?

OpenStudy (anonymous):

\[\tan^{-1} (\tan (4*\pi)/3)\]

OpenStudy (anonymous):

the reason it is not \[\frac{4\pi}{3}\] is that \[\frac{4\pi}{3}\] is larger than \[\frac{\pi}{2}\] and thefore not in the range of arctangent

OpenStudy (anonymous):

tangent is periodic with period pi, so subtract pi from \[\frac{4\pi}{3}\] so get \[\frac{\pi}{3}\]